OpenStudy (anonymous):

Solve: -4(2x - 1) = -6(x + 2) - 2

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

-4(2x-1)= -6(x+2)-2 -8x+4=-6x-12-2 combine the like terms -8x+6x=-12-2-4 -2x=-18

OpenStudy (anonymous):